OpenStudy (anonymous):

A system of equations is shown below: 5x + 2y = 3 (equation 1) 2x – 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof

OpenStudy (anonymous):

Answer choices: Show that the solution to the system of equations 7x – y = 4 and 2x – 3y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 2x + 5y = 3 and 3x – 2y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 9x + 4y = 5 and 7x – y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations –4x + 9y = 5 and 2x – 3y = 1 is the same as the solution to the given system of equations

OpenStudy (anonymous):

@Orion1213 can you help me with this one too?

OpenStudy (anonymous):

... i'm analyzing the problem... will get some coffee first...

OpenStudy (math&ing001):

(equation1) is replaced by [ (equation1) + 1*(equation2) ]. Choice 1 is correct.

OpenStudy (anonymous):

how @math&ing001 ?

OpenStudy (anonymous):

i agree with @math&ing001 ...

OpenStudy (anonymous):

Choice 1 is the same as solving the given equations...

OpenStudy (math&ing001):

Simply sum up the two equations to get the first one. (5x + 2y = 3) +(2x – 3y = 1)

OpenStudy (anonymous):

... combine like terms such as adding all x's and y's and constants...

OpenStudy (anonymous):

you'll get 7x-y=4...

OpenStudy (anonymous):

Thanks so much, you guys. I understand it now :) @Orion1213 @math&ing001

OpenStudy (math&ing001):

Welcome !